OpenStudy (diamond77108):

plz help I will medal −5(2x−6)+8x=22

OpenStudy (eatme):

what does x=?

OpenStudy (eatme):

if i give u the answer u must fan me

OpenStudy (diamond77108):

ok

OpenStudy (greison14):

then u have to fan her as well

OpenStudy (diamond77108):

who

OpenStudy (greison14):

the user eatme

OpenStudy (diamond77108):

i did

OpenStudy (eatme):

sorry wifi on crack right now

OpenStudy (diamond77108):

lmao

OpenStudy (eatme):

is there any answer choices

OpenStudy (kittiwitti1):

You want to solve for \(x\)?

OpenStudy (bamboo888):

x=4

OpenStudy (eatme):

wth

OpenStudy (eatme):

nah ur wrong

OpenStudy (kittiwitti1):

No direct answering please, at least explain part of it. @Bamboo888

OpenStudy (bamboo888):

I am, just wait

OpenStudy (eatme):

double check ur work

OpenStudy (kittiwitti1):

\[−5(2x−6)+8x=22\]Distributive property: \((-5(2x)-(-5)(6))+8x=22\) which leads to \((-10x-(-30))+8x=22\)

OpenStudy (kittiwitti1):

\[-10x+30+8x=22\]Add like terms and isolate \(x\), what did you get?

OpenStudy (eatme):

yup so far sooooooooo gooood

OpenStudy (kittiwitti1):

@diamond77108 ?

OpenStudy (eatme):

-2

OpenStudy (kittiwitti1):

Please don't give direct answers...

OpenStudy (eatme):

srry

OpenStudy (bamboo888):

It is 4 but I have to go, enter it on a calculator, but i have to go sorry, can someone help me explain?

OpenStudy (diamond77108):

-2x+30=22 -2x=-8

OpenStudy (kittiwitti1):

Yeah, I explained @Bamboo888

OpenStudy (diamond77108):

x=4

OpenStudy (kittiwitti1):

Yep

OpenStudy (diamond77108):

thx

OpenStudy (kittiwitti1):

No problem :-)

OpenStudy (bamboo888):

Thank @kittiwitti1